# A representational framework for learning and encoding structurally enriched trajectories in complex agent environments

### v1

@@ 3 Methodology
A key distinction of our framework is the ability to handle shared nodes that may appear in multiple trajectories or episodes. These shared nodes serve as bridges between episodes, allowing the emergence of cross-episode patterns. For example:
  • 1.
    An object, such as a ball, might be used across different episodes but in varying contexts (e.g., being pushed to a goal in one task and placed in a bucket in another).
  • 2.
    Similarly, an action like push could have different effects depending on the affordances and states in multiple episodes.
@@
We first define the ontology of our framework then describe the process of constructing Structurally Enriched Trajectories (SETs) as heterogeneous graphs, ensuring they encode multi-level relationships between objects, interactions, states, and affordances. We then detail the methodology for learning SET embeddings, which can be used for downstream tasks, including integration into reinforcement learning pipelines.
@@ 3.1.3 Structurally Enriched Trajectories (SETs)
is the set of edges, encoding temporal transitions $(s_{t},\mathfrak{\hat{f}}_{t},s_{t+1})$ and structural relationships such as object interactions $(o_{i},o_{j})$.
@@
3.1.4 Hierarchical Memory Structure
@@
3.1.5 Structured Data Source: CREATE Environment
The constraints of our framework required an environment with multiple tools and manipulable elements, allowing different solution strategies within and across tasks. To ensure the capture of structured dependencies over time, tasks needed to involve multi-step execution, reinforcing hierarchical trajectory learning rather than isolated decision-making. Additionally, task goals had to vary across episodes, compelling agents to adapt their strategies dynamically instead of relying on memorized action sequences. We employ Construction, Reuse, and Extension of Actions through Transfer and Exploration (CREATE) environment (Jain et al.,, 2020), a diverse and challenging platform specifically designed to study action generalisation in reinforcement learning. It provides a range of tasks that involve dynamic interactions with various objects and tools, such as pushing a ball to a designated area, placing it inside a container, or navigating obstacles. It also offers a rich setting to demonstrate the effectiveness of our approach, with tools to explore generalisation by varying task setups across episodes, forcing agents to adapt their strategies and reuse knowledge from past experiences.
@@ 3.2.1 SET Definition Through Random Exploration
A SET is labelled as success if the agent reaches the goal; otherwise, it is labelled as failure. The SET of an episode $SET(E)$ is labeled as follows:
@@
3.2.2 Hierarchical Memory Structure
To effectively model SETs, we structured episodic data into hierarchical heterogeneous graphs, where the feature extraction functions $\Phi:V\to\mathbb{R}^{d}$ (see Section 3.1.3) assigned high-dimensional embeddings to nodes representing objects, interactions, and states. The nodes and their different relations are organised into hierarchical heterogeneous graphs with three distinct levels of abstraction: (1) a base layer for objects and interactions, where objects are extracted using the SAM model (Kirillov et al.,, 2023) and interaction dynamics are encoded via a ConvLSTM; (2) a mid-layer for states and affordances, capturing the transition dynamics and relationships between interactions and their effects; and (3) a trajectory layer for task-level abstractions, representing the cumulative impact of interactions and sequences (see Fig. 2). This hierarchical organisation is stored in a graph database and is not simply a stratification of layers but a structured representation where lower levels influence higher-level abstractions. Objects and actions influence affordances, which in turn shape state transitions and contribute to the formation of SETs. This dependency across levels enables agents to reason over task structure, capturing both local dynamics and high-level task generalisation in complex environments. Formally, the hierarchical memory graph G representing a SET consists of:
Figure 1: This illustration represents the hierarchical structure of a SET, capturing relationships between different node types in task representation. At the lowest level, interactions (light-filled circular nodes) depend on objects (darker circular nodes). Moving upward, affordances (smaller darker nodes) emerge from interactions, which then influence states (highlighted in red with distinct borders) which in turn influence the next affordance. At the highest level, the SET node connects to the structure through temporal state dependencies. The edges indicate relationships such as object dependencies, contributions, and effects. Varying line styles help differentiate these relationships. This hierarchical structure showcases how task execution is represented by linking objects, interactions, affordances, and states.
@@ 3.2.2 Hierarchical Memory Structure
The lowest level of the hierarchy captures individual objects (Fig. 3) and interactions representations (Fig. 4), which form the fundamental units in the environment (Fig. 5).
Figure 3: Base-level representations: Object-level representations extracted at the base layer using the SAM object detection model. These representations capture object features in the environment and are stored in a graph database.
Figure 4: Base-level representations: Interaction-level representations extracted at the base layer using a pre-trained ConvLSTM action embedder. These embeddings capture the dynamics of object interactions and are stored in the graph database.
@@
Figure 5: Representation of two objects interacting. Interactions are characterized not only by their execution but also by the relationships they share with the objects involved. On the right, the different parameters of an Interaction are shown such as the type, and the vector representation. Representation obtained from the Neo4j (Lal,, 2015) graph database interface.
@@
    If $\text{If }\max_{i}\,\text{Similarity}(o_{\text{new}},o_{i})<\tau,\text{ add }o_{\text{new}}\text{ as a new node.}$. (8)
@@
    Let $\mathcal{F}_{a}=\{f_{1},f_{2},\dots,f_{n}\mid\text{all frames associated with interaction }a\}$ represent the set of frames associated with a specific interaction $a$. For a given frame $x_{a}\in\mathcal{F}_{a}$, we denote $x_{p}\in\mathcal{F}_{a}$ as a positive sample (a frame associated with the same interaction $a$) and $x_{n}\notin\mathcal{F}_{a}$ as a negative sample (a frame associated with a different interaction).
@@
    Figure 6: The scatter plot visualises the second object in interactions categorized as Fixed (orange, right side of the plot) or Bouncy (blue, left side of the plot), with each symbol representing a unique object (e.g. triangle, hexagon etc). The type of symbol along with their color indicates specific objects (e.g., Bouncy_Triangle, Fixed_Hexagon). The y-axis represents unique object IDs, while the x-axis categorizes objects into Bouncy or Fixed.
@@
Figure 7: Illustration of the hierarchical structure in SETLE as extracted from the Neo4j graph database (Lal,, 2015). The second level of the hierarchy consists of states (red, from the middle node the first set of connections) connected by affordances (green, the outer-most nodes), illustrating the transitions between states via affordances.
@@
The hierarchical graph structure obtained illustrated in Figure 1, captures the relationships between different levels of abstraction within an episode. The visualisation shows a diverse set of nodes, each representing a unique component of the episode, including objects, states, interactions, and affordances, linked through various relationships.
@@ 3.3 SETLE: SET encoder
A graph encoder is trained with triplet loss to ensure that episodes with similar trajectories cluster together, distinguishing successful task executions from unsuccessful ones.
3.3.1 Collecting data for training
We collect SETs and label them to create a dataset for training, by following these steps:
  1. 1.
    Random exploration: We begin with an exploration phase where the agent performs multiple tasks within the CREATE environment. During this phase, a random policy agent operates in the environment until it reaches the goal state. For tractability, we have limited the number of steps the agent can take in an environment to 9, if it has not reached a goal state by then, the SET is labelled as unsuccessful. We repeat the process until sufficient successful trajectories have been collected for training.
  2. 2.
@@ 3.3.2 Graph Encoder Architecture Based on Heterogeneous Co-contrastive Learning (HeCo)
A graph encoder is a neural architecture that transforms graph-structured data into meaningful embeddings, preserving the relationships and dependencies between entities. Unlike standard feature extractors, graph encoders operate over relational data, using the graph topology to learn structured representations.
@@
    The selected metapaths, SET-St-Obj-Inter-Aff and SET-St-Aff-St, are designed to capture high-order semantic relationships within each episode in the RL environment.
@@
        This metapath is valuable for modelling temporal dependencies in SETs, as it reveals the sequences of states the agent encounters. Understanding these transitions is critical for building a representation that reflects the path to the goal, rather than just the individual states, thus helping the agent generalise across different task setups that share similar transition dynamics.
@@
4 Experiments
This section evaluates the effectiveness of SETLE’s hierarchical graph encoder in encoding task-specific information. The evaluation is structured into two key sets of experiments: margin sensitivity analysis and ablation studies. In the first set of experiments (S1), we assessed how different margin values in the triplet loss function affect the quality of learnt embeddings. By varying the margin parameter, we analysed its influence on the clustering of success and failure SETs in the embedding space. This experiment provided insights into the optimal margin settings that enhanced separability while maintaining robust intra-cluster cohesion.
@@ 4.1 S1: SET Embedding Analysis: Margin and Loss Function Impact
The evaluation focused on specific parameters including:
@@
      Hybrid Triplet Loss: Combines triplet loss with a cross-view contrastive loss. The contrastive component aligns embeddings from the schema and meta-path views, while the triplet component enhances task-based separability.

### v2

@@ 3 Methodology
A key distinction of our framework is the ability to handle shared nodes that may appear in multiple trajectories or episodes. These shared nodes serve as bridges between episodes, allowing the emergence of cross-episode patterns. For example an object, such as a ball, might be used across different episodes but in varying contexts (e.g., being pushed to a goal in one task and placed in a bucket in another). Similarly, an action like push could have different effects depending on the affordances and states in multiple episodes.
@@
We first define the ontology of our framework in Subsection 3.1,. then describe the process of constructing Structurally Enriched Trajectories (SETs) as heterogeneous graphs Subsection 3.2, ensuring they encode multi-level relationships between objects, interactions, states, and affordances. We next detail the methodology for learning SET embeddings in Subsection 3.3, which can be used for downstream tasks, including integration into reinforcement learning pipelines as seen in Section 5.
@@ 3.1.3 Structurally Enriched Trajectories (SETs)
is the set of edges, encoding temporal transitions $(s_{t},\mathfrak{\hat{f}}_{t},s_{t+1})$ and structural rrelationships, such as object interactions $(o_{i},o_{j})$.
@@
3.1.4 Ontology of the Hierarchical Memory
@@
This formal definition of the memory as a hierarchical, heterogeneous graph provides the structural foundation for our learning framework. The specific architecture used to process this structure and the mechanism by which knowledge is propagated across its shared nodes are detailed in Section 3.3.
3.1.5 Structured Data Sources: CREATE and MiniGrid
The constraints of our framework required environments that support structured, multi-step interactions and diverse tasks. To test our approach across different types of complexity, we selected two distinct platforms.
Our primary environment is Construction, Reuse, and Extension of Actions through Transfer and Exploration (CREATE) (Jain et al.,, 2020). CREATE is a diverse and challenging platform designed to study action generalisation, providing tasks that involve dynamic interactions with various objects and tools in a continuous physics-based simulation. It allows us to explore generalisation by varying task setups, forcing agents to adapt strategies and reuse knowledge from past experiences.
To complement CREATE’s physical complexity, we also employ the MiniGrid benchmark suite. MiniGrid offers a simplified, discrete setting with minimalist 2D grid worlds composed of symbolic entities like keys and doors. It features a limited, discrete action space and agent-centric partial observability. This environment serves to test SETLE’s capacity for abstraction and multi-step planning in a domain that emphasizes logical and planning complexity rather than perceptual richness.
@@ 3.2.1 SET Definition Through Random Exploration
A SET is labelled as success if the agent reaches the goal; otherwise, it is labelled as failure. The SET of an episode $SET(E)$ is labelled as follows:
@@
3.2.2 Building the Hierarchical Memory Graph
To effectively model SETs, we structured episodic data into hierarchical heterogeneous graphs, where the feature extraction functions $\Phi:V\to\mathbb{R}^{d}$ (see Section 3.1.3) assigned high-dimensional embeddings to nodes representing objects, interactions, and states. The nodes and their different relations are organised into hierarchical heterogeneous graphs with three distinct levels of abstraction: (1) a base layer for objects and interactions, where objects are extracted using the SAM model (Kirillov et al.,, 2023) and interaction dynamics are encoded via a ConvLSTM; (2) a mid-layer for states and affordances, capturing the transition dynamics and relationships between interactions and their effects; and (3) a trajectory layer for task-level abstractions, representing the cumulative impact of interactions and sequences (see Fig. 2). The selection of these specific models was based on a comparative analysis to ensure robustness and generalisability, as explained in our component analysis in Section 4.2. This hierarchical organisation is stored in a graph database and is not simply a stratification of layers but a structured representation where lower levels influence higher-level abstractions. Objects and actions influence affordances, which in turn shape state transitions and contribute to the formation of SETs. This dependency across levels enables agents to reason over task structure, capturing both local dynamics and high-level task generalisation in complex environments. Formally, the hierarchical memory graph G representing a SET consists of:
Figure 1: An example of the Hierarchical Structure of a SET Graph, illustrating Multi-Level Relational Dependencies. At the lowest level, interactions (light-filled circular nodes) depend on objects (darker circular nodes). Moving upward, affordances (smaller darker nodes) emerge from interactions, which then influence states (highlighted in red with distinct borders) which in turn influence the next affordance. At the highest level, the SET node connects to the structure through temporal state dependencies. The edges indicate relationships such as object dependencies, contributions, and effects. Varying line styles help differentiate these relationships. This hierarchical structure showcases how task execution is represented by linking objects, interactions, affordances, and states.
@@ 3.2.2 Building the Hierarchical Memory Graph
The lowest level of the hierarchy captures individual objects (Fig. 3 nodes in orange) and interactions representations (Fig. 3 node in pink), which form the fundamental units in the environment (Fig. 3).
@@
Figure 3: Representation of two objects interacting. Interactions are characterized not only by their execution , but also by the relationships they share with the objects involved. On the right, the different parameters of an Interaction are shown such as the type, and the vector representation. Representation obtained from the Neo4j (Lal,, 2015) graph database interface.
@@
    $\text{If }\max_{i}\,\text{Similarity}(o_{\text{new}},o_{i})<\tau,\text{ add }o_{\text{new}}\text{ as a new node.}$ (8)
@@
[…]
@@
    Figure 4: Visualisation of the interaction encoder’s ability to learn semantically meaningful representations. This scatter plot shows K-means clusters of interaction embeddings, categorized by the physical properties of the object involved (’Bouncy’ vs. ’Fixed’). The clear separation between the two groups demonstrates that the encoder has successfully learned to distinguish between these two functionally critical classes of interactions from visual data alone, a key prerequisite for building the SET hierarchy.
@@
Figure 5: Illustration of the hierarchical structure in SETLE representing the second level of the hierarchy that consists of states (red, from the middle node the first set of connections) connected by affordances (green, the outer-most nodes), illustrating the transitions between states via affordances.
@@
The hierarchical graph structure obtained, illustrated in Figure 1, captures the relationships between different levels of abstraction within an episode. The visualisation shows a diverse set of nodes, each representing a unique component of the episode, including objects, states, interactions, and affordances, linked through various relationships.
@@ 3.3 SETLE: SET encoder
A graph encoder is trained with triplet loss to ensure that episodes with similar trajectories cluster together, distinguishing successful task executions from unsuccessful ones (Fig.6).
Figure 6: Overview of the SETLE encoder training pipeline. Stage 1: A triplet of SETs (an anchor, a positive, and a negative sample) is used as input. Stage 2: All three subgraphs are processed by an encoder with shared parameters (f(x; $\theta$)), which uses the HeCo architecture to generate embeddings from both the Network Schema and Meta-path views. These views are regularized by an internal cross-view contrastive loss. Stage 3: The final embeddings are used to compute the Triplet Loss, which pulls the anchor and positive samples together in the latent space while pushing the negative sample away.
@@ 3.3.1 Graph Encoder Architecture Based on Heterogeneous Co-contrastive Learning (HeCo)
To operate on the heterogeneous graphs and leverage the meta-path relationships formally defined in our Hierarchical Memory Structure (Section 3.1.4), we employ a specialized graph encoder. Specifically, we adapt the Heterogeneous Co-contrastive Learning (HeCo) architecture that transforms graph-structured data into meaningful embeddings, preserving the relationships and dependencies between entities. Unlike standard feature extractors, graph encoders operate over relational data, using the graph topology to learn structured representations.
@@
    The selected meta-paths, SET-St-Obj-Inter-Aff and SET-St-Aff-St, are designed to capture high-order semantic relationships within each episode in the RL environment.
@@
        This meta-path is valuable for modelling temporal dependencies in SETs, as it reveals the sequences of states the agent encounters. Understanding these transitions is critical for building a representation that reflects the path to the goal, rather than just the individual states, thus helping the agent generalise across different task setups that share similar transition dynamics.
@@
3.3.2 Implicit Knowledge Propagation via Shared Embeddings
As established in Section 3.1.4, shared nodes serve as the ”bridges” for knowledge transfer. This section details the implicit learning mechanism that makes this transfer possible during the encoder’s training. A key feature of SETLE’s design is the implicit propagation of knowledge across episodic subgraphs during the training of the encoder. While there is no direct message-passing operating between different subgraphs during a single forward pass, knowledge is transferred through the shared, updatable embeddings of common nodes during the iterative training process.
The process, is implemented as follows:
  • 1.
    Independent Subgraph Encoding: At each training step, a triplet of self-contained episode subgraphs (an anchor, a positive, and a negative sample) is loaded. The features and graph structure for each are prepared independently before being passed to the model, ensuring message passing is confined within the boundaries of each individual trajectory.
  • 2.
    Shared Node Embeddings: A single, globally-stored feature embedding table is maintained within the HeCo model for each node type. This updatable table is referenced via unique node IDs every time a concept (e.g., the ‘red ball‘ object) appears, ensuring the same vector representation is used across all episodes.
  • 3.
    Update via Backpropagation with Gradient Accumulation: During training, gradients from the hybrid loss function, calculated on the output of the episode triplet, flow back via the loss.backward() call. These gradients are accumulated over several steps (e.g., 20 steps in our experiments). The model’s parameters, including the shared embedding layers, are then updated synchronously. This use of gradient accumulation allows the model to aggregate signals from multiple triplets before updating the shared embeddings, leading to more stable learning.
  • 4.
    Implicit Knowledge Transfer: The model’s parameters are persistent across the entire training loop. Therefore, an update to a shared node’s embedding from a triplet processed at step $i$ is immediately available when that node appears in another triplet at a subsequent step $i+k$. This iterative refinement is how learning from one experience implicitly informs the processing of another. To use an analogy, the shared nodes function like recurring characters in a narrative; their state (embedding) is shaped by the events of one chapter (an episode triplet) and they carry that updated state with them into subsequent chapters.
3.3.3 Collecting data for training
To populate the Hierarchical Memory Structure ($\mathcal{M}$) with the individual SET subgraphs ($\mathcal{G}_{i}$) needed for the training process described above, and specifically to generate the anchor, positive, and negative samples required by our hybrid triplet loss, we first generate a corpus of SETs from agent interactions. These trajectories are then labelled by their task outcome (success or failure) to create a dataset for training, following these steps:
  1. 1.
    Random exploration: We begin with an exploration phase where the agent performs multiple tasks within the CREATE and MiniGrid environment. During this phase, a random policy agent operates in the environment until it reaches the goal state. For tractability, we have limited the number of steps the agent can take in an environment to 9, if it has not reached a goal state by then, the SET is labelled as unsuccessful. We repeat the process until sufficient successful trajectories have been collected for training.
  2. 2.
4 Multi-Environment Experiments
This section evaluates the effectiveness of SETLE’s hierarchical graph encoder in encoding task-specific information across environments with different complexity profiles. The evaluation is structured around two key experimental domains: the CREATE environment, which tests perceptual and physical complexity, and the MiniGrid environment, which tests logical and planning complexity.
While the CREATE environment may appear visually simple, it provides a unique advantage as a testbed. It maintains a low computational overhead, yet incorporates complex, continuous physics dynamics —such as a ball being propelled by a fan or bouncing off different surfaces—that are often absent in grid-world settings. This unique combination makes it an ideal environment for studying foundational principles like tool use and affordance learning, without requiring the extensive processing power of high-fidelity simulators. In contrast, MiniGrid offers a simplified, discrete setting that tests SETLE’s capacity for abstraction and multi-step planning under partial observability.
The evaluation is structured into two key sets of experiments: margin sensitivity analysis and ablation studies. In the first set of experiments (S1), we assessed how different margin values in the triplet loss function affect the quality of learnt embeddings. By varying the margin parameter, we analysed its influence on the clustering of success and failure SETs in the embedding space. This experiment provided insights into the optimal margin settings that enhanced separability while maintaining robust intra-cluster cohesion.
@@ 4 Multi-Environment Experiments
In the following subsections, we first present the detailed results from the CREATE environment, which has a richer hierarchical structure including interaction dynamics. We then present the results from MiniGrid, where the graph structure is simplified, as interaction-level dynamics cannot be extracted meaningfully. This allows us to test SETLE’s capacity for abstraction under reduced representational complexity.
4.1 Experimental Setup and Implementation Details
This section provides a comprehensive overview of the model architecture, training configuration, and hyperparameters used in our experiments.
Our model architecture is based on HeCo (Heterogeneous Co-contrastive Learning) with a GraphSAGE encoder. The encoder uses a 3-layer architecture with 64 hidden units per layer, producing node embeddings with 64 hidden dimensions. The architecture is configured to process 2 meta-paths, as defined in Section 3.3.1, and the attention mechanism employs 8 heads with 64-dimensional key/value vectors. To ensure reproducibility, all random seeds for PyTorch, NumPy, and Python’s random were fixed at the beginning of each run. We used the PyTorch default initialization methods: Xavier/Glorot for linear layers and He for convolutional layers.
For training, we used the Adam optimizer with a learning rate of 0.001 and a weight decay (L2 regularization) coefficient of 1e-4. To simulate a larger batch size and stabilize training, gradients were accumulated over 20 steps before each parameter update. The encoder was trained for up to 200 epochs, with early stopping based on validation loss. The training objective was a hybrid loss function that combines a triplet objective with a cross-view contrastive loss.
To ensure the robustness of our findings for the encoder, performance was assessed over 8 different runs. For the downstream reinforcement learning experiments presented in Section 5, each agent configuration was trained and evaluated across 3 different random seeds. The data for the encoder training was sourced from agent interactions and divided using a standard 80-20 train/validation split. The model was trained on the training set with early stopping based on performance on the validation set to prevent overfitting.
4.2 Component Analysis and Design Justification
Beyond evaluating the complete framework, it is important to justify the design choices for its critical prerequisite components. Rather than performing traditional ablations on these foundational parts, this section analyses their contribution by comparing alternative approaches and highlighting the framework’s modularity across different environments.
ConvLSTM Encoder for Interaction Dynamics
As introduced in Section 3.2.2, the ConvLSTM encoder is designed to capture rich, temporal interaction dynamics from visual input, which is essential in physics-based environments like CREATE. Our multi-environment evaluation serves as a natural experiment for this component’s role. In the symbolic MiniGrid environment, where there are no complex visual dynamics to model, the ConvLSTM and its corresponding ‘ActionRepr‘ nodes are deliberately omitted from the graph structure. The framework’s success in both domains demonstrates that the ConvLSTM is a powerful but modular component: its contribution is to enable the processing of complex perceptual dynamics when they exist, but the core SETLE framework is robust enough to function without it in logically-driven domains.
Object Discovery: Foundational Models for Robust Perception
A prerequisite for our framework is a robust object discovery module that can generalize across different visual domains without environment-specific retraining. In our initial research, we evaluated traditional object-centric models, specifically Slot Attention. While effective on its training data, we found that it failed to generalize to the visual styles of our RL environments, producing unstable segmentations even after fine-tuning. In contrast, a foundational model—the Segment Anything Model (SAM), provided excellent, zero-shot object segmentation across all tested domains without any modification. This design choice was critical for ensuring the robustness and generalizability of the entire SETLE pipeline, which would otherwise be bottlenecked by a brittle, domain-specific perception module.
@@ 4.3 S1: SET Embedding Analysis: Margin and Loss Function Impact
The evaluation focused on specific pparameters, including:
@@
      Hybrid Triplet Loss: Combines ta triplet loss with a cross-view contrastive loss. The contrastive component aligns embeddings from the schema and meta-path views, while the triplet component enhances task-based separability.